THE INTERMOLECULAR STRUCTURE OF THE PHENOL DIMER AND THE BENZONITRILE DIMER IN THE S_{0} AND S_{1} STATE DETERMINED BY HIGH RESOLUTION ROTATIONALLY RESOLVED ELEC-TRONIC SPECTROSCOPY

\underline{W. LEO MEERTS}, IVO KALKMAN, GIEL BERDEN^{a} AND MARLOES VAN BEEK, Molecularand Biophysics Group, Institute for Molecules and Materials, Radboud University Nijmegen, P.O. Box 9010, NL-6500 GL Nijmegen, The Netherlands; \underline{M. SCHMITT}, MARCEL BÖHM, CHRISTIAN RATZER, DANIEL KRÜGLER^{b}, SWEN SIEGERT AND KARL KLEINERMANNS, Heinrich-Heine-Universität, Institut für Physikalische Chemie, D-40225 Düsseldorf, Germany..

The rotationally resolved UV spectra of the electronic origins of five isotopomers of the phenol dimer and the most abundant isotopomer of the benzonitrile dimer have been measured. The complex spectra are analyzed using a genetic algorithm based fitting strategy. For the phenol dimer we determined from the inertial parameters the intermolecular geometry parameters for the electronic states S_{0} and S_{1} and compared them to the results of ab initio calculations. In the electronic ground state S_{0} a larger hydrogen bond length than in the ab initio calculations is found and a smaller tilt angle of the aromatic rings, showing a more pronounced dispersion interaction. In the electronically excited state S_{1} the hydrogen bond length decreases, as has been found for other hydrogen bonded clusters of phenol and the two aromatic rings are tilted less toward each other. For the benzinitrile dimer we found that the electronic ground state has a C_{2h} symmetric structure in which the permanent dipole moments of the benzonitrile monomers are aligned anti-parallel. The orientation of the transition dipole moment could be shown to be parallel to the orientation in the monomer moiety. The distance between the two monomer moieties decreases slightly upon electronic excitation and the symmetry of the benzonitrile dimer changes from C_{2h} in the electronic ground state to C_{s} in the electronically excited state. This break of symmetry is probably caused by the local excitation of only one benzonitrile moiety in the cluster due to the weak electronic coupling between the cluster moieties.

aPresent address: FOM Institute for Plasma Physics Rijnhuizen, Edisonbaan 14, 3439 MN Nieuwegein, The Netherlands

bPresent address: Bruker Daltonik GmbH, D-28359 Bremen, Germany